# Scalable non-separable spatio-temporal Gaussian process models for large-scale short-term weather prediction

### v1

@@ Abstract
Monitoring daily weather fields is critical for climate science, agriculture, and environmental planning, yet fully probabilistic spatio-temporal models become computationally prohibitive at continental scale. We present a case study on short-term forecasting of daily maximum temperature and precipitation across the conterminous United States using novel scalable spatio-temporal Gaussian process methodology. Building on three approximation families — inducing-point methods (FITC), Vecchia approximations, and a hybrid Vecchia–inducing-point full-scale approach (VIF) — we introduce three extensions that address key bottlenecks in large space-time settings: (i) a scalable correlation-based neighbor selection strategy for Vecchia approximations with point-referenced data, enabling accurate conditioning under complex dependence structures, (ii) a space-time kMeans++ inducing-point selection algorithm, and (iii) GPU-accelerated implementations of computationally expensive operations, including matrix operations and neighbor searches. Using both synthetic experiments and a large NOAA station dataset containing approximately 1.7 million space-time observations, we analyze the models with respect to predictive performance, parameter estimation, and computational efficiency. Our results demonstrate that scalable Gaussian process models can yield accurate continental-scale forecasts while remaining computationally feasible, offering practical tools for weather applications.
@@ 1 Introduction
Reliable short-term weather forecasts are essential for various applications including crop management, water allocation, flood preparedness, and utility planning. High-resolution weather observations from dense meteorological networks and remote sensing platforms have created new opportunities for statistical modeling of atmosphere-land data. Daily temperature and precipitation fields exhibit substantial variability across spatial scales, are strongly modulated by topography and large-scale circulation, and evolve with seasonal cycles. Forecasting these fields, therefore, requires accounting for joint space-time dependence rather than treating locations or time points independently. Operational weather forecasts are typically based on numerical weather prediction (NWP) models (Bauer et al., 2015), which solve nonlinear differential equations governing atmospheric motion and quantify uncertainty using ensemble simulation. While highly successful, NWP systems demand specialized high-performance computing resources and are not always available at the temporal resolution, geographic domain, or computational budget needed for regional planning, ecological monitoring, or exploratory scientific analysis. In many such settings, statistical models that learn directly from observations offer a valuable complement. Gaussian Process (GP) models provide a flexible, fully probabilistic, and data-driven approach to representing complex space-time structure. In particular, non-separable covariance functions, such as those of Gneiting (2002), allow for spatial correlation to vary with temporal lag and temporal correlation to depend on spatial separation, capturing interactions between space and time that are essential for realistic weather dynamics. These capabilities explain their widespread use in environmental and climate statistics, including weather and short-term spatio-temporal modeling (Cressie and Wikle, 2011; Porcu et al., 2021).
@@ 4.1 Vecchia approximation
$\boldsymbol{b}\stackrel{{\scriptstyle\text{approx}}}{{\sim}}\mathcal{N}(0,{\mathbf{\Sigma}}_{s}),\qquad{\mathbf{\Sigma}}_{s}=\boldsymbol{B}^{-1}\boldsymbol{D}\boldsymbol{B}^{-T},$
with a sparse precision matrix ${\mathbf{\Sigma}}_{s}^{-1}=\boldsymbol{B}^{T}\boldsymbol{D}^{-1}\boldsymbol{B}$. Here, $\boldsymbol{B}\in\mathbb{R}^{n\times n}$ is a unit lower-triangular matrix whose nonzero off-diagonal entries in row $i$ occur only in the columns indexed by $N(i)$ and are given by $\boldsymbol{A}_{i}\;=\;\mathbf{\Sigma}_{i,N(i)}\,{\mathbf{\Sigma}}_{N(i)}^{-1}$, where $\mathbf{\Sigma}_{i,N(i)}$ denotes the cross-covariances between $b((\boldsymbol{s}_{i},t_{i}))$ and its Vecchia neighbors $\{b((\boldsymbol{s}_{k},t_{k})):k\in N(i)\}$, and ${\mathbf{\Sigma}}_{N(i)}$ is the covariance submatrix indexed by $N(i)$. The diagonal matrix $\boldsymbol{D}=\mathrm{diag}(D_{1},\ldots,D_{n})\in\mathbb{R}^{n\times n}$ is given by $D_{i}\;=\;{\mathbf{\Sigma}}_{i,i}-\boldsymbol{A}_{i}\,\mathbf{\Sigma}_{i,N(i)}^{\mathrm{T}}$.
@@ 4.2 Low-rank inducing point approximations
$\boldsymbol{b}\stackrel{{\scriptstyle\text{approx}}}{{\sim}}\mathcal{N}(0,{\mathbf{\Sigma}}_{l}),\qquad{\mathbf{\Sigma}}_{l}=\mathbf{\Sigma}_{mn}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mn}+\mathrm{diag}\!\big({\mathbf{\Sigma}}-\mathbf{\Sigma}_{mn}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mn}\big).$
@@ 4.3 The Vecchia-inducing-point full-scale (VIF) approximation
VIF approximations introduced in Gyger et al. (2025b) combine the strengths of low-rank and Vecchia methods through a full-scale decomposition. The latent process is decomposed as $b((\boldsymbol{s},t))=b_{l}((\boldsymbol{s},t))+b_{s}((\boldsymbol{s},t))$, where $b_{l}$ is a low-rank predictive process capturing large-scale structure and $b_{s}$ is a residual process capturing local dependence not explained by $b_{l}$. The covariance of the low-rank component is $\mathbf{\Sigma}_{l}=\mathbf{\Sigma}_{mn}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mn}$, while the residual covariance is $\mathrm{Cov}(\boldsymbol{b}_{s})={\mathbf{\Sigma}}-\mathbf{\Sigma}_{mn}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mn}$. This residual covariance is dense, and it is approximated using a Vecchia approximation $\mathrm{Cov}(\boldsymbol{b}_{s})^{-1}\approx{\mathbf{\Sigma}}_{s}^{-1}=\boldsymbol{B}^{\top}\boldsymbol{D}^{-1}\boldsymbol{B}$, where the Vecchia coefficients now use residual covariances adjusted for the low-rank component. Specifically, $D_{i}={\mathbf{\Sigma}}_{i,i}-\mathbf{\Sigma}_{mi}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mi}-\boldsymbol{A}_{i}\big(\mathbf{\Sigma}_{i,N(i)}^{\mathrm{T}}-\mathbf{\Sigma}_{mN(i)}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mi}\big)$, and $\boldsymbol{A}_{i}=\big(\mathbf{\Sigma}_{i,N(i)}-\mathbf{\Sigma}_{mi}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mN(i)}\big)\big({\mathbf{\Sigma}}_{N(i)}-\mathbf{\Sigma}_{mN(i)}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mN(i)}\big)^{-1}$. The resulting full-scale approximation is $\mathbf{\Sigma}\approx\mathbf{\Sigma}_{l}+{\mathbf{\Sigma}}_{s}$.
@@ 4.4.1 Correlation-based Vecchia neighbor selection
Nearest-neighbor search with respect to both correlation metrics can be implemented efficiently using cover trees (Beygelzimer et al., 2006). We use an extension of the cover tree algorithm to enable efficient correlation-based neighbor selection for general, non-separable space-time covariance functions. In contrast to standard constructions that randomly select candidate nodes during tree building, we insert points sequentially according to their index order, following the procedure outlined in Algorithm 2 in Appendix B. This modification restricts candidate neighbors to observations with smaller indices, which is consistent with the Vecchia factorization and simplifies the neighbor search, leading to improved runtime performance, see Algorithm 3 in Appendix B for more details. Because covariance parameters evolve during likelihood optimization, neighbor sets are updated at optimization iterations indexed by powers of two and once more upon convergence. If an update alters the likelihood, optimization is resumed until convergence is achieved.
@@ 5 Simulated experiments
We first compare the predictive and inferential performance of the above-introduced GP approximations using simulated experiments. Specifically, we evaluate Vecchia approximations with either Euclidean or correlation-based nearest-neighbor selection, both using $m_{v}=30$ neighbors. For low-rank inducing-point methods, we consider FITC approximations with inducing points selected via both the standard kMeans++ and the proposed sts-kMeans++ algorithm, in both cases using $m=500$ inducing points. Finally, we evaluate the hybrid VIF approximation, which combines correlation-based Vecchia neighbor selection with sts-kMeans++-selected inducing points, using $m_{v}=30$ and $m=500$. All experiments are conducted on a system equipped with an AMD EPYC 7763 processor using 16 CPU threads, 512 GB of system memory, and an NVIDIA L40S GPU with 48 GB of VRAM. GPU-based computations refer to executions that utilize the L40S GPU in addition to the CPU, whereas CPU-based computations are performed exclusively on the AMD EPYC 7763 processor. All methods are implemented using GPBoost, and the complete R and Python source code and datasets required to reproduce the experiments are publicly available at https://github.com/TimGyger/SpaceTimeGPApprox.
@@ 6 Application to short-term weather prediction
In the following, we apply the proposed methodology to generating short-term forecasts of daily maximum temperature and precipitation across the conterminous United States using the datasets and models described in Sections 2 and 3, respectively. The observational network comprises 3,057 stations with highly heterogeneous spatial coverage (see Figures 5 and 6 in Appendix A). Such variability directly motivates scalable space-time modeling, as accurate prediction must transfer information across regions with widely differing station density. The dataset contains 608 consecutive days from 2024-01-01 to 2025-08-31. Forecasts are produced using an expanding-window framework. All observations from 2024 serve as the initial training set. This initial training dataset contains more than one million space-time observations. Beginning on January 1, 2025, we sequentially generate forecasts on every day for daily temperatures and rainfall with lead times of five and three days, respectively. The covariance parameters are re-estimated at the end of each month which allows for adapting to temporal non-stationarity (e.g., seasonal shifts in variability and correlation structure). Compared to a daily re-estimation, this reduces computation times while retaining operational realism. We use the same computational environment including GPU acceleration and the same number of Vecchia neighbors and inducing points as in Section 5.
To evaluate prediction accuracy, a random subset of approximately 25% of the stations is withheld entirely during training. This spatio-temporal test data design ensures that every model is evaluated not only for temporal forecasting at observed sites but also for its ability to transfer information to previously unmonitored locations. In practice, forecast users often require predictions in areas with sparse observational histories, at newly installed sensors, or in regions without any monitoring infrastructure (e.g., agricultural zones, remote mountain terrain, coastal transition areas). The Python code used to run all experiments is publicly accessible at https://github.com/TimGyger/SpaceTimeGPApprox.
@@ 6.1 Short-term prediction of daily maximum temperature
Table 7 in Appendix E reports the overall RMSE and CRPS aggregated across the full five-day horizon for pure temporal, and spatio-temporal, together with average computation times. The VIF approximation provides the highest prediction accuracy, albeit at a higher training cost than the Vecchia methods. The FITC approximation is computationally inexpensive but substantially less accurate. Across all scalable approximations, forecast generation is inexpensive: computing predictive means and variances requires only a few tens of seconds. Thus, the dominant computational expense arises from parameter estimation rather than daily forecasting. Even for the most accurate method (VIF), a full re-estimation takes on the order of one hour, while the Vecchia variants complete training within minutes. These runtimes remain practical even without GPU acceleration. Even on standard CPU hardware, monthly parameter re-estimation can be performed offline (e.g., overnight) after which operational 5-day forecasts can be issued almost instantly. Consequently, scalable GP models can be deployed in routine forecasting workflows without specialized hardware, while still delivering high-accuracy, uncertainty-aware predictions at continental scale.
@@ 6.2 Short-term prediction of precipitation
For precipitation, we use the covariance structure (2) with the Matérn smoothness fixed to $\nu=0.5$. This corresponds to an exponential covariance, producing a rougher latent process that more realistically captures the sharp spatial and temporal variability characteristic of rainfall. Estimating the smoothness parameter using only CPU on subsets of the data gives $\nu\approx 0.5$. Daily precipitation exhibits a mixed distribution with a discrete component (occurrence) and a continuous component (amount). To account for this, we use the zero-censored power-transformed normal (ZC-PTN) likelihood introduced in (3) in Section 3.2. A Laplace approximation is used for parameter estimation and prediction. To reduce the computational complexity, we employ the iterative and stochastic techniques of Kündig and Sigrist (2025) and Gyger et al. (2025b), which replace explicit matrix factorizations with iterative methods that only require matrix-vector operations. To evaluate the predictive accuracy, we use separate metrics for occurrences and intensities. Because skill in precipitation forecasting deteriorates more rapidly with lead time than for temperature, we focus on one- to three-day-ahead forecasts, an operationally relevant horizon at which meaningful predictability can still be obtained. We also examined up to five-day-ahead forecasts; however, accuracy remained largely similar or only slightly degraded. For rainfall occurrence, we use the Brier score (BS) and the log score (LS), which assess both sharpness and calibration of the implied occurrence probabilities. Conditional on rainfall occurrence, positive-amount predictions are evaluated using the mean absolute error (MAE), which is interpretable on the original scale and robust to the heavy right tail of precipitation amounts. To assess probabilistic predictions, we additionally consider the continuous ranked probability score (CRPS), computed by sampling from the ZC-PTN predictive distributions. All GP-based methods are compared with a fixed-effects linear regression model without a GP and a simple persistence benchmark based on the previous day’s precipitation occurrence and amount.

### v2

@@ Abstract
Monitoring daily weather fields is critical for climate science, agriculture, and environmental planning, yet fully probabilistic spatio-temporal models become computationally prohibitive at continental scale. We present a case study on short-term forecasting of daily maximum temperature and precipitation across the conterminous United States using novel scalable spatio-temporal Gaussian process methodology. Building on three approximation families — inducing-point methods (FITC), Vecchia approximations, and a hybrid Vecchia–inducing-point full-scale approach (VIF) — we introduce three extensions that address key bottlenecks in large space-time settings: (i) a scalable correlation-based neighbor selection strategy for Vecchia approximations with point-referenced data, enabling accurate conditioning under complex dependence structures, (ii) a space-time kMeans++ inducing-point selection algorithm, and (iii) GPU-accelerated implementations of computationally expensive operations, including matrix operations and neighbor searches. Using both synthetic experiments and a large NOAA station dataset containing more than one million space-time observations, we analyze the models with respect to predictive performance, parameter estimation, and computational efficiency. Our results demonstrate that scalable Gaussian process models can yield accurate continental-scale forecasts while remaining computationally feasible, offering practical tools for weather applications.
@@ 1 Introduction
Reliable short-term weather forecasts are essential for various applications including crop management, water allocation, hydrological modeling, water-dependent risk assessment, and utility planning. High-resolution weather observations from dense meteorological networks and remote sensing platforms have created new opportunities for statistical modeling of atmosphere-land data. Daily temperature and precipitation fields exhibit substantial variability across spatial scales, are strongly modulated by topography and large-scale circulation, and evolve with seasonal cycles. Forecasting these fields, therefore, requires accounting for joint space-time dependence rather than treating locations or time points independently. Operational weather forecasts are typically based on numerical weather prediction (NWP) models (Bauer et al., 2015), which solve nonlinear differential equations governing atmospheric motion and quantify uncertainty using ensemble simulation. While highly successful, NWP systems demand specialized high-performance computing resources and are not always available at the temporal resolution, geographic domain, or computational budget needed for regional planning, ecological monitoring, or exploratory scientific analysis. In many such settings, statistical models that learn directly from observations offer a valuable complement. Gaussian Process (GP) models provide a flexible, fully probabilistic, and data-driven approach to representing complex space-time structure. In particular, non-separable covariance functions, such as those of Gneiting (2002), allow for spatial correlation to vary with temporal lag and temporal correlation to depend on spatial separation, capturing interactions between space and time that are essential for realistic weather dynamics. These capabilities explain their widespread use in environmental and climate statistics, including weather and short-term spatio-temporal modeling (Cressie and Wikle, 2011; Porcu et al., 2021).
@@ 4.1 Vecchia approximation
$\boldsymbol{b}\stackrel{{\scriptstyle\text{approx}}}{{\sim}}\mathcal{N}(0,{\mathbf{\Sigma}}^{\mathrm{s}}),\qquad{\mathbf{\Sigma}}^{\mathrm{s}}=\boldsymbol{B}^{-1}\boldsymbol{D}\boldsymbol{B}^{-T},$
with a sparse precision matrix $({\mathbf{\Sigma}}^{\mathrm{s}})^{-1}=\boldsymbol{B}^{T}\boldsymbol{D}^{-1}\boldsymbol{B}$. Here, $\boldsymbol{B}\in\mathbb{R}^{n\times n}$ is a unit lower-triangular matrix whose nonzero off-diagonal entries in row $i$ occur only in the columns indexed by $N(i)$ and are given by $\boldsymbol{A}_{i}\;=\;\mathbf{\Sigma}_{i,N(i)}\,{\mathbf{\Sigma}}_{N(i)}^{-1}$, where $\mathbf{\Sigma}_{i,N(i)}$ denotes the cross-covariances between $b((\boldsymbol{s}_{i},t_{i}))$ and its Vecchia neighbors $\{b((\boldsymbol{s}_{k},t_{k})):k\in N(i)\}$, and ${\mathbf{\Sigma}}_{N(i)}$ is the covariance submatrix indexed by $N(i)$. The diagonal matrix $\boldsymbol{D}=\mathrm{diag}(D_{1},\ldots,D_{n})\in\mathbb{R}^{n\times n}$ is given by $D_{i}\;=\;{\mathbf{\Sigma}}_{i,i}-\boldsymbol{A}_{i}\,\mathbf{\Sigma}_{i,N(i)}^{\mathrm{T}}$.
@@ 4.2 Low-rank inducing point approximations
$\boldsymbol{b}\stackrel{{\scriptstyle\text{approx}}}{{\sim}}\mathcal{N}(0,{\mathbf{\Sigma}}^{\mathrm{l}}),\qquad{\mathbf{\Sigma}}^{\mathrm{l}}=\mathbf{\Sigma}_{mn}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mn}+\mathrm{diag}\!\big({\mathbf{\Sigma}}-\mathbf{\Sigma}_{mn}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mn}\big).$
@@ 4.3 The Vecchia-inducing-point full-scale (VIF) approximation
VIF approximations introduced in Gyger et al. (2025b) combine the strengths of low-rank and Vecchia methods through a full-scale decomposition. The latent process is decomposed as $b((\boldsymbol{s},t))=b_{\mathrm{l}}((\boldsymbol{s},t))+b_{\mathrm{s}}((\boldsymbol{s},t))$, where $b_{\mathrm{l}}$ is a low-rank predictive process capturing large-scale structure and $b_{\mathrm{s}}$ is a residual process capturing local dependence not explained by $b_{\mathrm{l}}$. The covariance of the low-rank component is $\mathbf{\Sigma}^{\mathrm{l}}=\mathbf{\Sigma}_{mn}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mn}$, while the residual covariance is $\mathrm{Cov}(\boldsymbol{b}_{s})={\mathbf{\Sigma}}-\mathbf{\Sigma}_{mn}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mn}$. This residual covariance is dense, and it is approximated using a Vecchia approximation $\mathrm{Cov}(\boldsymbol{b}_{s})^{-1}\approx{(\mathbf{\Sigma}}^{\mathrm{s}})^{-1}=\boldsymbol{B}^{\top}\boldsymbol{D}^{-1}\boldsymbol{B}$, where the Vecchia coefficients now use residual covariances adjusted for the low-rank component. Specifically, $D_{i}={\mathbf{\Sigma}}_{i,i}-\mathbf{\Sigma}_{mi}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mi}-\boldsymbol{A}_{i}\big(\mathbf{\Sigma}_{i,N(i)}^{\mathrm{T}}-\mathbf{\Sigma}_{mN(i)}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mi}\big)$, and $\boldsymbol{A}_{i}=\big(\mathbf{\Sigma}_{i,N(i)}-\mathbf{\Sigma}_{mi}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mN(i)}\big)\big({\mathbf{\Sigma}}_{N(i)}-\mathbf{\Sigma}_{mN(i)}^{\mathrm{T}}\mathbf{\Sigma}_{m}^{-1}\mathbf{\Sigma}_{mN(i)}\big)^{-1}$. The resulting full-scale approximation is $\mathbf{\Sigma}\approx\mathbf{\Sigma}^{\mathrm{l}}+{\mathbf{\Sigma}}^{\mathrm{s}}$.
@@ 4.4.1 Correlation-based Vecchia neighbor selection
Nearest-neighbor search with respect to both correlation metrics can be implemented efficiently using cover trees (Beygelzimer et al., 2006). We use an extension of the cover tree algorithm to enable efficient correlation-based neighbor selection for general, non-separable space-time covariance functions. In contrast to standard constructions that randomly select candidate nodes during tree building, we insert points sequentially according to their index order, following the procedure outlined in Algorithm 2 in Appendix B. This modification restricts candidate neighbors to observations with smaller indices, which is consistent with the Vecchia factorization and simplifies the neighbor search, leading to improved runtime performance, see Algorithm 3 in Appendix B for more details. Because covariance parameters evolve during optimization, neighbor sets are updated at iterations indexed by powers of two and once more upon convergence. If an update alters the likelihood, the optimization is resumed until convergence.
@@ 5 Simulated experiments
We first compare the predictive and inferential performance of the above-introduced GP approximations using simulated experiments. Specifically, we evaluate Vecchia approximations with either Euclidean or correlation-based nearest-neighbor selection, both using $m_{v}=30$ neighbors. For low-rank inducing-point methods, we consider FITC approximations with inducing points selected via both the standard kMeans++ and the proposed sts-kMeans++ algorithm, in both cases using $m=500$ inducing points. Finally, we evaluate the hybrid VIF approximation, which combines correlation-based Vecchia neighbor selection with sts-kMeans++-selected inducing points, using $m_{v}=30$ and $m=500$. All experiments are conducted on a system equipped with an AMD EPYC 7763 processor using 16 CPU threads, 512 GB of system memory, and an NVIDIA L40S GPU with 48 GB of VRAM. GPU-based computations refer to executions that utilize the L40S GPU in addition to the CPU, whereas CPU-based computations are performed exclusively on the AMD EPYC 7763 processor. All methods presented in this article are implemented in the GPBoost library (https://github.com/fabsig/GPBoost), and we use version 1.6.6 to perform the simulated experiments and the real-wold application. The code and datasets required to reproduce all experiments are publicly available at https://github.com/TimGyger/SpaceTimeGPApprox.
@@ 6 Application to short-term weather prediction
In the following, we apply the proposed methodology to generating short-term forecasts of daily maximum temperature and precipitation across the conterminous United States using the datasets and models described in Sections 2 and 3. The observational network comprises 3,057 stations with highly heterogeneous spatial coverage (see Figures 5 and 6 in Appendix A). Such variability directly motivates scalable space-time modeling, as accurate prediction must transfer information across regions with widely differing station density. The dataset contains 608 consecutive days from 2024-01-01 to 2025-08-31. Forecasts are produced using expanding-window training datasets with observations from the year 2024 serving as initial training set. To evaluate the prediction accuracy, a random subset of approximately 25% of the stations is withheld entirely during training. This spatio-temporal test data design ensures that every model is evaluated not only for temporal forecasting at observed sites but also for its ability to transfer information to previously unmonitored locations. In practice, forecast users often require predictions in areas with sparse observational histories, at newly installed sensors, or in regions without any monitoring infrastructure (e.g., agricultural zones, remote mountain terrain, coastal transition areas). Beginning on January 1, 2025, we sequentially generate forecasts on every day for daily temperatures and rainfall with lead times of five and three days, respectively. The covariance parameters are re-estimated at the end of each month which allows for adapting to temporal non-stationarity (e.g., shifts in variability and correlation structure). Compared to a daily re-estimation, this reduces computation times while retaining operational realism. We use the same computational environment including GPU acceleration and the same number of Vecchia neighbors and inducing points as in Section 5. The training datasets contain between approximately $800,000$ and $1.3$ million space-time points.
@@ 6.1 Short-term prediction of daily maximum temperature
Table 7 in Appendix E reports the average RMSE and CRPS across the full five-day horizon and average runtimes. Recall that the training datasets contain between approximately $800,000$ and $1.3$ million space-time points. The VIF approximation provides the highest prediction accuracy, albeit at a higher training cost than the Vecchia methods. The FITC approximation is computationally inexpensive but substantially less accurate. Across all scalable approximations, forecast generation is inexpensive: computing predictive means and variances requires only a few tens of seconds. Thus, the dominant computational expense arises from parameter estimation rather than daily forecasting. Even for the most accurate method (VIF), a full re-estimation takes on the order of one hour, while the Vecchia variants complete training within minutes. These runtimes remain practical even without GPU acceleration. Even on standard CPU hardware, monthly parameter re-estimation can be performed offline (e.g., overnight) after which operational 5-day forecasts can be issued almost instantly. Consequently, scalable GP models can be deployed in routine forecasting workflows without specialized hardware, while still delivering high-accuracy, uncertainty-aware predictions at continental scale.
@@ 6.2 Short-term prediction of precipitation
For precipitation, we use the covariance in (2) with the Matérn smoothness fixed to $\nu=0.5$. This corresponds to an exponential covariance, producing a rougher latent process that more realistically captures the sharp spatial and temporal variability characteristic of rainfall. Estimating the smoothness parameter using only CPU on subsets of the data gives $\nu\approx 0.5$. Daily precipitation exhibits a mixed distribution with a discrete component (occurrence) and a continuous component (amount). To account for this, we use the zero-censored power-transformed normal (ZC-PTN) likelihood introduced in (3) in Section 3.2. A Laplace approximation is used for parameter estimation and prediction. To reduce the computational complexity, we use the iterative methods of Kündig and Sigrist (2025) and Gyger et al. (2025b), which replace explicit matrix factorizations with iterative methods that only require matrix-vector operations. To evaluate the predictive accuracy, we use separate metrics for occurrences and intensities. Because skill in precipitation forecasting deteriorates more rapidly with lead time than for temperature, we focus on one- to three-day-ahead forecasts, an operationally relevant horizon at which meaningful predictability can still be obtained. We also examined up to five-day-ahead forecasts; however, accuracy remained similar or only slightly degraded. For rainfall occurrence, we use the Brier score (BS) and the log score (LS), which assess both sharpness and calibration of the implied occurrence probabilities. Conditional on rainfall occurrence, positive-amount predictions are evaluated using the mean absolute error (MAE), which is robust to the heavy right tail of precipitation amounts. To assess probabilistic predictions, we additionally consider the CRPS, computed by sampling from the ZC-PTN predictive distributions. For comparison, we include a fixed-effects linear regression model without a GP and a simple persistence benchmark based on the previous day’s precipitation occurrence and amount.